The question is solve:
log(x)3^12=24
I am not quite sure how to solve this. I moved the equation into
x^24=3^12
x^24=531441
but I am stuck now and I am not sure if I even am on the right track.

Answer by Alan3354(69443)

log(x)3^12=24
I am not quite sure how to solve this. I moved the equation into
x^24=3^12
x^2 = 3
x = sqrt(3)
=============
x^24=531441